Shift report 20010918 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Sep 18 2001 - 00:15:39 EDT

  • Next message: Apache: "Shift report 20010918 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Sep 18 00:15:39 2001, Shift: 16:00-24:00
    Supervisor: Y. K. Lee, with: 
    Log book#8 pages 149 - 151
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine was cogged at top energy with 56 bunches and
       200 M Herz for 40 minutes until there was a fire alarm
       in 1008B. It gave an improved Z-vertex distribution of
       sigma=20 cm. After vacuume trouble, machine was cogged
       at 22:15 and still running the storage.
    2. Run 4777, one half hour run with Trigger 1:77 K events.
       Run 4780, since 22:15, 258 K events and still running.
    3. ZDC ADC Right module 1(Blue) was dead. Hiro essentially
       jiggled the connectors and it came back.
       H1 TD Bottom 33 is dead.
       TOFWall Cal ADC TOP3 has wide pedestal.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Sep 18 2001 - 00:16:11 EDT